# Quantum Error Correction near the Coding Theoretical Bound

### v2

@@ A.2 Finite field extension to $(H_{\Gamma},H_{\Delta})$
[…]
@@
The matrix in equation (9) can be transformed into its reduced row echelon form using elementary row operations without requiring division. By randomly selecting a random solution of this system, we obtain $\underline{\lambda}$ followed by $\underline{\gamma}$ and $H_{\Gamma}$. Furthermore, treating the non-zero entries of $H_{\Delta}$ as variables, the linear equation $H_{\Gamma}H_{\Delta}^{\mathsf{T}}=0$ is solved to determine the non-zero entries of $H_{\Delta}$.
@@ A.3 Conversion to $(H_{X},H_{Z})$
[…]
@@
[…]
@@
From the properties of the companion matrix (Appendix B), it can be verified that $(H_{X},H_{Z})$ are orthogonal.
@@
Appendix B Properties of the companion matrix
@@
[…]
@@
[…]
@@
Appendix C Details of decoding methods
C.1 Decoding algorithm
Let $M=PJ$, $N=PL$, and $n=eN$. Using the matrices $H_{X},H_{Z}\in\mathbb{F}_{2}^{eM\times eN}$ and $H_{\Gamma},H_{\Delta}\in\mathbb{F}_{q}^{M\times N}$ provided in the previous sections, we describe the method for performing error correction. The decoding process involves estimating the noise vectors $\underline{x},\underline{z}\in\mathbb{F}_{2}^{eN}$ from the syndromes $\underline{s}=H_{Z}\underline{x}$ and $\underline{t}=H_{X}\underline{z}$, where the matrices $H_{X}$ and $H_{Z}$ are constructed from the non-binary matrices $H_{\Gamma}$ and $H_{\Delta}$ as described in the previous section.
@@ C.1 Decoding algorithm
$R=1-H(p_{D},p_{D}/3,p_{D}/3,p_{D}/3)=1+(1-3p_{D})\log_{2}(1-3p_{D})+3p_{D}\log_{2}(p_{D}),$ (41)
where $H(\cdot)$ is the entropy function. The probability of occurrence for $\underline{x},\underline{z}\in\mathbb{F}_{2}^{eN}$ in this channel is given by the following expression:
@@
$p(x,z)$ =$=\begin{cases}1-p_{D},&(x,z)=(0,0),\\ \frac{p_{D}}{3},&(x,z)=(0,1),(1,0),(1,1),\end{cases}$ (44)
@@
The posterior probability of $\underline{x},\underline{z}$ given the syndromes $\underline{s},\underline{t}$ can be sparsely factorized as follows:
$p(\underline{x},\underline{z}|\underline{s},\underline{t})$ $\propto\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{Z})}_{ij}x_{j}=s_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{X})}_{ij}z_{j}=t_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{j=1}^{N}p\left(x_{j},z_{j}\right)\Bigr{)},$ (45)
where $(H_{X})_{ij}$, $(H_{Z})_{ij}$ denotes the $(i,j)$-th $(e\times e)$ submatrix of $H_{X},H_{Z}$, respectively. We denote $\mathbbm{1}[\cdot]$ as 1 if the inside the brackets is true, and 0 otherwise. For each $i$, note that there are $L$ values of $j$ for which $(H_{X})_{ij}$ and $(H_{Z})_{ij}$ are non-zero, respectively.
@@
The SP algorithm that marginalizes this posterior probability (45), specifically, is an algorithm that updates the following eight types of messages:
$\mu^{X}_{ij}(x_{j}),\mu^{Z}_{ij}(z_{j}),\nu^{X}_{ji}(x_{j}),\nu^{Z}_{ji}(z_{j}),\lambda^{X}_{j}(x_{j}),\lambda^{Z}_{j}(z_{j}),\kappa^{X}_{j}(x_{j}),\kappa^{Z}_{j}(z_{j}).$
All messages are initialized to a uniform distribution over $x_{j}$ and $z_{j}$. Each message is updated using the following update equations. After each update, the messages are normalized to ensure they form probability distributions.
$\nu_{ij}^{X}(x_{j})$ $=\sum_{(x_{j^{\prime}}):{j^{\prime}\in\partial_{X}(i)}\setminus j}\mathbbm{1}\Bigl{[}\sum_{j^{\prime}\in\partial_{X}(i)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{Z})}_{ij^{\prime}}x_{j^{\prime}}=s_{i}\Bigr{]}\prod_{j^{\prime}\in\partial_{X}(i)\setminus j}\mu_{j^{\prime}i}^{X}(x_{j^{\prime}}),$ (46)
$\nu_{ij}^{Z}(z_{j})$ $=\sum_{(z_{j^{\prime}}):{j^{\prime}\in\partial_{Z}(i)}\setminus j}\mathbbm{1}\Bigl{[}\sum_{j^{\prime}\in\partial_{Z}(i)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{X})}_{ij^{\prime}}z_{j^{\prime}}=t_{i}\Bigr{]}\prod_{j^{\prime}\in\partial_{Z}(i)\setminus j}\mu_{j^{\prime}i}^{Z}(z_{j^{\prime}}),$ (47)
$\lambda_{j}^{X}(x_{j})$ $=\prod_{i\in\partial_{X}(j)}\nu_{ij}^{X}(x_{j}),$ (48)
$\lambda_{j}^{Z}(z_{j})$ $=\prod_{i\in\partial_{Z}(j)}\nu_{ij}^{Z}(z_{j}),$ (49)
@@
$\mu^{X}_{ji}(x_{j})$ $=\kappa^{X}_{j}(x_{j})\prod_{i^{\prime}\in\partial_{X}(j)\setminus i}\nu_{i^{\prime}j}^{X}(x_{j}),$ (52)
$\mu^{Z}_{ji}(z_{j})$ $=\kappa^{Z}_{j}(z_{j})\prod_{i^{\prime}\in\partial_{Z}(j)\setminus i}\nu_{i^{\prime}j}^{Z}(z_{j}),$ (53)
where $\partial_{X}(i)$ and $\partial_{Z}(i)$ are the set of column index $j$ such that $(H_{X})_{ij}\neq O$ and $(H_{Z})_{ij}\neq O$, respectively. Similarly, $\partial_{X}(j)$ and $\partial_{Z}(j)$ are the set of row index $i$ such that $(H_{X})_{ij}\neq O$ and $(H_{Z})_{ij}\neq O$, respectively. It holds that #$\#\partial_{X}(i)=\#\partial_{Z}(i)=J$ and #$\#\partial_{X}(j)=\#\partial_{Z}(j)=L$.
@@
$\hat{x}_{j}=\operatornamewithlimits{argmax}_{x_{j}}\prod_{i\in\partial_{X}(j)}\nu_{ij}^{X}(x_{j}),\quad\hat{z}_{j}=\operatornamewithlimits{argmax}_{z_{j}}\prod_{i\in\partial_{Z}(j)}\nu_{ij}^{Z}(z_{j})$ (54)
@@
$\sum_{j\in\partial_{X}(i)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{Z})}_{ij}\hat{x}_{j}=s_{i},\quad\sum_{j\in\partial_{Z}(i)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{X})}_{ij}\hat{z}_{j}=t_{i},$
then $(\hat{x}_{j})$ and $(\hat{z}_{j})$ are accepted as the estimated noise, and the algorithm terminates. If the syndrome are not satisfied, a decoding failure is declared.
The messages (46) and (47) can be computed using FFT [18], requiring $O(Lq\log q)$ operations per message. This message computation can also be implemented as a decoding method for LDPC codes over a finite field. The detail is provided in Appendix C.2. On the other hand the computation for messages (52) and (53) require $q^{2}$ operations per message. Other messages computation require only $O(Jq)$ operations per message.
@@
C.2 Sum-product algorithm over finite field
We show $p(\underline{x},\underline{z}|\underline{s},\underline{t})$ can also be expressed using elements of a finite field as follows:
$p(\underline{x},\underline{z}|\underline{s},\underline{t})\propto\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta_{ij}}\xi_{j}=\sigma_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma_{ij}}\zeta_{j}=\tau_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{j=1}^{N}p\left(x_{j},z_{j}\right)\Bigr{)}.$ (55)
For $s_{i},t_{i},x_{j},z_{j}\in\mathbb{F}_{2}^{e}$, we define $\sigma_{i},\tau_{i},\xi_{j},\zeta_{j}\in\mathbb{F}_{q}$ are such that $\underline{w}(\sigma_{i})=s_{i},\underline{v}(\tau_{i})=t_{i},\underline{w}(\xi_{j})=x_{j}$ and $\underline{v}(\zeta_{j})=z_{j}$ hold. Note that $\underline{v}(\cdot)$ and $\underline{w}(\cdot)$ were defined in B.
From Bayes’ rule and the fact that the observed syndromes $\underline{s},\underline{t}$ can be treated as constants, we have:
$p(\underline{x},\underline{z}|\underline{s},\underline{t})$ $=p(\underline{s},\underline{t}|\underline{x},\underline{z})p(\underline{x},\underline{z})/p(\underline{s},\underline{t})$ (56)
$\propto p(\underline{s},\underline{t}|\underline{x},\underline{z})p(\underline{x},\underline{z}).$ (57)
Next, using the chain rule and the fact that the syndromes are determined by the noise, we obtain:
$p(\underline{s},\underline{t}|\underline{x},\underline{z})$ $=p(\underline{s},\underline{t}|\underline{x},\underline{z})$ (58)
$=p(\underline{s}|\underline{t},\underline{x},\underline{z})p(\underline{t}|\underline{x},\underline{z})$ (59)
$=p(\underline{s}|\underline{x})p(\underline{t}|\underline{z}).$ (60)
From the fact that the noise and syndromes are determined by the parity-check equations and the properties of the parity-check matrices, we have:
$p(\underline{t}|\underline{z})$ $=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}(H_{X})_{ij}z_{j}=t_{i}\bigr{]}$ (61)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A(\gamma_{ij})z_{j}=t_{i}\bigr{]}$ (62)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A(\gamma_{ij})\underline{v}(\zeta_{j})=\underline{v}(\tau_{i})\bigr{]}$ (63)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}\delta_{ij}\zeta_{j}=\tau_{i}\bigr{]}.$ (64)
For similar reasons, we obtain the following. Note that the companion matrix is given by $A$ and the vector representation is given by $\underline{v}(\cdot)$:
$p(\underline{s}|\underline{x})$ $=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}(H_{Z})_{ij}x_{j}=s_{i}\bigr{]}$ (65)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A^{\mathsf{T}}(\delta_{ij})x_{j}=s_{i}\bigr{]}$ (66)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A^{\mathsf{T}}(\delta_{ij})\underline{w}(\xi_{j})=\underline{w}(\sigma_{i})\bigr{]}$ (67)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}\delta_{ij}\xi_{j}=\sigma_{i}\bigr{]}.$ (68)

### v3

@@ 5.1 Finite field extension to $(H_{\Gamma},H_{\Delta})$
First, let us determine the non-zero entries $(\gamma_{ij})$ of $H_{\Gamma}$. Consider a one-dimensional array of length $2LP$, denoted as $\underline{\gamma}=(\gamma_{0},\gamma_{1},\ldots,\gamma_{2LP-1})$, which satisfies the following conditions: for $0\leq i<P$, $\gamma_{ij}=\gamma_{j}$, and for $P\leq i<2P$, $\gamma_{ij}=\gamma_{j+PL}$. Assume that $\mathbb{F}_{q}$ contains a primitive element $\alpha$. Thus, $\mathbb{F}_{q}$ can be written as $\mathbb{F}_{q}=\{0,\alpha^{0},\alpha^{1},\ldots,\allowbreak\alpha^{q-2}\}$. We assume that the condition (a) is satisfied. Let us consider the condition for the existence of a non-zero $H_{\Gamma}$ that satisfies the orthogonality equation $H_{\Gamma}H_{\Delta}^{\mathsf{T}}=0$, treating $H_{\Gamma}$ as a variable. Denote $\underline{\lambda}\overset{\text{\small def}}{=}(\log_{\alpha}\gamma_{0},\ldots,\log_{\alpha}\gamma_{2PL-1})^{\mathsf{T}}\in\mathbb{Z}_{q-1}^{2PL}$. When condition (b) holds, all-non-zero solution $(\gamma_{ij})$ of the equation $H_{\Gamma}H_{\Delta}^{\mathsf{T}}=0$ exist if and only if the following system of linear congruences over $\mathbb{Z}_{q-1}$ with respect to $\underline{\lambda}$:
@@
The matrix in equation (10) can be transformed into its blackuced row echelon form using elementary row operations without requiring division. By randomly selecting a random solution of this system, we obtain $\underline{\lambda}$ followed by $\underline{\gamma}$ and $H_{\Gamma}$. Furthermore, treating the non-zero entries of $H_{\Delta}$ as variables, the linear equation $H_{\Gamma}H_{\Delta}^{\mathsf{T}}=0$ is solved to determine the non-zero entries of $H_{\Delta}$.
@@ 5.2 Conversion to $(H_{X},H_{Z})$
$A(\alpha)$ $\overset{\text{\small def}}{=}\begin{bmatrix}0&0&0&0&a_{0}\\ 1&0&0&0&a_{1}\\ 0&1&0&0&a_{2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&1&a_{e-1}\end{bmatrix}$. (11)
@@
$A(0)$ $\overset{\text{\small def}}{=}O,$ (12)
$A(\alpha^{l})$ $\overset{\text{\small def}}{=}A(\alpha)^{l}\quad\text{for }l=0,\ldots,q-2.$ (13)
@@
From the properties of the companion matrix in the following subsection, it can be verified that $(H_{X},H_{Z})$ are orthogonal.
@@
5.3 Properties of the companion matrix
@@
$A^{\mathsf{T}}(0)$ $\overset{\text{\small def}}{=}O,$ (27)
$A^{\mathsf{T}}(\alpha^{i})$ $\overset{\text{\small def}}{=}(A(\alpha)^{\mathsf{T}})^{i}\quad\text{for }i=0,\ldots,q-2.$ (28)
@@
Define $\underline{w}(0)\overset{\text{\small def}}{=}(0,\ldots,0)\in\mathbb{F}_{2}^{e}$. Let $\underline{w}(\alpha^{i})\in\mathbb{F}_{2}^{e}$ be defined as the first column of $A^{\mathsf{T}}(\alpha^{i})$. This correspondence establishes a group isomorphism between $\mathbb{F}_{q}$ and $\mathbb{F}_{2}^{e}$: $\underline{w}(\gamma_{1}+\gamma_{2})=\underline{w}(\gamma_{1})+\underline{w}(\gamma_{2})$. The following properties hold:
@@
5.4 Details of decoding methods
[…]
@@ 5.4 Details of decoding methods
$R=1-H_{2}(p_{D})-p_{D}\log_{2}(3),$ (42)
where $H_{2}(\cdot)$ is the binary entropy function. The probability of occurrence for $\underline{x},\underline{z}\in\mathbb{F}_{2}^{eN}$ in this channel is given by the following expression:
@@
$p(x,z)$ $=\begin{cases}1-p_{D},&(x,z)=(0,0),\\ \frac{p_{D}}{3},&(x,z)=(0,1),(1,0),(1,1),\end{cases}$ (45)
@@
Using Bayes’ rule, the posterior probability of $\underline{x},\underline{z}$ given the syndromes $\underline{s},\underline{t}$ can be sparsely factorized as follows:
$p(\underline{x},\underline{z}|\underline{s},\underline{t})$ ${\color[rgb]{0,0,0}=p(\underline{s},\underline{t}|\underline{x},\underline{z})p(\underline{x},\underline{z})/p(\underline{s},\underline{t})}$ (46)
${\color[rgb]{0,0,0}\propto p(\underline{s}|\underline{x})p(\underline{t}|\underline{z})p(\underline{x},\underline{z})}$ (47)
$=\Bigl(\prod_{i=1}^{M}\mathbbm{1}\bigl[\sum_{j=1}^{N}(H_{Z})_{ij}x_{j}=s_{i}\bigr]\Bigr)\Bigl(\prod_{i=1}^{M}\mathbbm{1}\bigl[\sum_{j=1}^{N}(H_{X})_{ij}z_{j}=t_{i}\bigr]\Bigr)\Bigl(\prod_{j=1}^{N}p\left(x_{j},z_{j}\right)\Bigr),$ (48)
where $(H_{X})_{ij}$, $(H_{Z})_{ij}$ denotes the $(i,j)$-th $(e\times e)$ submatrix of $H_{X},H_{Z}$, respectively. For each $i$, note that there are $L$ values of $j$ for which $(H_{X})_{ij}$ and $(H_{Z})_{ij}$ are non-zero, respectively. We denote $\mathbbm{1}[\cdot]$ as 1 if the inside the brackets is true, and 0 otherwise. This expression is well-suited for representing the probability distribution from which each $e$-bit noise vector $x_{i},z_{i}$ arises.
@@
The SP algorithm that marginalizes the posterior probability (48) proceeds by iteratively updating the following four types of messages:
$\mu^{X}_{ij}(x_{j}),\quad\nu^{X}_{ji}(x_{j}),\quad\lambda^{X}_{j}(x_{j}),\quad\kappa^{X}_{j}(x_{j})$
for all $(i,j)$ such that $(H_{X})_{ij}\neq O$. Each message is defined as a real-valued function over binary vectors of length $e$, or equivalently, as a real-valued array of size $2^{e}$. Here, following standard conventions in SP algorithms, we allow a slight abuse of notation by using $x_{j}$ and $z_{j}$ to refer both to the random variables and to the corresponding variable nodes over which the messages are defined. The SP algorithm operates by iteratively updating these messages.
In the terminology of SP algorithms, $\mu^{X}_{ij}(x_{j})$ denotes the message passed from the variable node $x_{j}$ to the factor node corresponding to the $i$th parity-check equation, while $\nu^{X}_{ji}(x_{j})$ is the message in the reverse direction. Similarly, $\lambda^{X}_{j}(x_{j})$ represents the message from the variable node $x_{j}$ to the factor node corresponding to the joint prior $p(x_{j},z_{j})$, and $\kappa^{X}_{j}(x_{j})$ is the message in the reverse direction. The messages on the $Z$ side, $\mu^{Z}_{ij}(z_{j}),\ \nu^{Z}_{ji}(z_{j}),\ \lambda^{Z}_{j}(z_{j}),\ \kappa^{Z}_{j}(z_{j})$, are defined analogously.
All messages are initialized to a uniform distribution over $x_{j}$ and $z_{j}$. For example, $\mu_{ij}^{X}\left(x_{j}\right)$ is initialized to $1/2^{e}$ for all values of $x_{j}$. Each message is updated using the following update equations.
$\nu_{ij}^{X}(x_{j})$ $=\sum_{({x}_{j^{\prime}}):{j^{\prime}\in\partial_{{\color[rgb]{0,0,0}Z}}(i)}\setminus j}\mathbbm{1}\Bigl[\sum_{j^{\prime}\in\partial_{{\color[rgb]{0,0,0}Z}}(i)}(H_{Z})_{ij^{\prime}}x_{j^{\prime}}=s_{i}\Bigr]\prod_{j^{\prime}\in\partial_{{\color[rgb]{0,0,0}Z}}(i)\setminus j}\mu_{j^{\prime}i}^{X}(x_{j^{\prime}}),$ (49)
$\nu_{ij}^{Z}(z_{j})$ $=\sum_{({z}_{j^{\prime}}):{j^{\prime}\in\partial_{{\color[rgb]{0,0,0}X}}(i)}\setminus j}\mathbbm{1}\Bigl[\sum_{j^{\prime}\in\partial_{{\color[rgb]{0,0,0}X}}(i)}(H_{X})_{ij^{\prime}}z_{j^{\prime}}=t_{i}\Bigr]\prod_{j^{\prime}\in\partial_{{\color[rgb]{0,0,0}X}}(i)\setminus j}\mu_{j^{\prime}i}^{Z}(z_{j^{\prime}}),$ (50)
$\lambda_{j}^{X}(x_{j})$ $=\prod_{i\in\partial_{{\color[rgb]{0,0,0}Z}}(j)}\nu_{ij}^{X}(x_{j}),$ (51)
$\lambda_{j}^{Z}(z_{j})$ $=\prod_{i\in\partial_{{\color[rgb]{0,0,0}X}}(j)}\nu_{ij}^{Z}(z_{j}),$ (52)
@@
$\mu^{X}_{ji}(x_{j})$ $=\kappa^{X}_{j}(x_{j})\prod_{i^{\prime}\in\partial_{{\color[rgb]{0,0,0}Z}}(j)\setminus i}\nu_{i^{\prime}j}^{X}(x_{j}),$ (55)
$\mu^{Z}_{ji}(z_{j})$ $=\kappa^{Z}_{j}(z_{j})\prod_{i^{\prime}\in\partial_{{\color[rgb]{0,0,0}X}}(j)\setminus i}\nu_{i^{\prime}j}^{Z}(z_{j}),$ (56)
where $\partial_{X}(i)$ and $\partial_{Z}(i)$ are the set of column index $j$ such that $(H_{X})_{ij}\neq O$ and $(H_{Z})_{ij}\neq O$, respectively. After each update, the messages are normalized to ensure they form probability distributions. Similarly, $\partial_{X}(j)$ and $\partial_{Z}(j)$ are the set of row index $i$ such that $(H_{X})_{ij}\neq O$ and $(H_{Z})_{ij}\neq O$, respectively. It holds that $\partial_{X}(i)=\#\partial_{Z}(i)={\color[rgb]{0,0,0}L}$ and $\partial_{X}(j)=\#\partial_{Z}(j)={\color[rgb]{0,0,0}J}$.
@@
$\hat{x}_{j}=\operatornamewithlimits{argmax}_{x_{j}}{\color[rgb]{0,0,0}\kappa_{j}^{X}(x_{j})}\prod_{i\in\partial_{{\color[rgb]{0,0,0}Z}}(j)}\nu_{ij}^{X}(x_{j}),\quad\hat{z}_{j}=\operatornamewithlimits{argmax}_{z_{j}}{\color[rgb]{0,0,0}\kappa_{j}^{Z}(z_{j})}\prod_{i\in\partial_{{\color[rgb]{0,0,0}X}}(j)}\nu_{ij}^{Z}(z_{j})$ (57)
@@
$\sum_{j\in\partial_{{\color[rgb]{0,0,0}Z}}(i)}(H_{Z})_{ij}\hat{x}_{j}=s_{i},\quad\sum_{j\in\partial_{{\color[rgb]{0,0,0}X}}(i)}(H_{X})_{ij}\hat{z}_{j}=t_{i},$
then $(\hat{x}_{j})$ and $(\hat{z}_{j})$ are accepted as the estimated noise, and the algorithm terminates. If the syndromes are not satisfied, a decoding failure is declablack.
The messages (49) and (50) can be computed using FFT [17], requiring $O(Lq\log q)$ operations per message. On the other hand the computation for messages (55) and (56) require $q^{2}$ operations per message. Other messages computation require only $O(Jq)$ operations per message.
@@
6 Data Availability
All data and figures supporting the main conclusions of this study are available from the corresponding author upon reasonable request. Please contact Kenta Kasai at kenta@ict.eng.isct.ac.jp.
7 Code Availability
The code used in this study is available from the corresponding author upon reasonable request. Please contact Kenta Kasai at kenta@ict.eng.isct.ac.jp.
8 Acknowledgements
Not applicable.